This time we add a caption by means of the table environment:

\begin{table}[h]          % ...h for here (rather than, e.g., top)
\begin{center}
\begin{tabular}{|l|l|l|}
\hline
No. & Name & Function \\
\hline
1.  & CONTACT & to define a contact \\
2.  & END     & to end the data input \\
\hline
\end{tabular}
\caption{This is a table caption.}
\end{center}
\end{table}

No. Name  Function
1. CONTACT to define a contact
2. END  to end the data input

Table 1: This is a table caption.